OpenStudy (anonymous):

What is the Least Common Denominator of the fraction 2 over x plus the fraction x over x plus 1 plus the fraction x plus 1 over 3 x ?

OpenStudy (anonymous):

Please help?

OpenStudy (anonymous):

it is really hard to read this in words, the equation editor below is your best bet

OpenStudy (anonymous):

my guess is \[\frac{2}{x}+\frac{x}{x+1}+\frac{1}{3x}\]

OpenStudy (anonymous):

of course i could be wrong

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

suppose \(x=7\) what is the least common multiple of \(7,8,21\)?

OpenStudy (anonymous):

No clue

OpenStudy (anonymous):

A pretty large number

OpenStudy (phi):

any idea what the common denominator is for 1/x and 1/(3x) ?

OpenStudy (anonymous):

Umm.

OpenStudy (anonymous):

Just x?

OpenStudy (anonymous):

I need this to finish my test Please hurry in helping

OpenStudy (phi):

if you have \[ \frac{1}{x} + \frac{1}{3x} \] you need the bottom (denominator) to be the same for both fractions I would multiply the first fraction by 3/3 so you get 3x in the bottom if you need to add a third fraction, multiply the denominators